Clopack Company manufactures one product that goes through one processing department called Mixing. All raw materials are introduced at the start of work in the Mixing Department. The company uses the weighted-average method of process costing. Its Work in Process T-account for the Mixing Department for June follows (all forthcoming questions pertain to June):

Work in Process—Mixing Department
June 1 balance

38,000

Completed and transferred
to Finished Goods
?
Materials 123,705
Direct labor 82,500
Overhead 100,000
June 30 balance ?

The June 1 work in process inventory consisted of 5,300 units with $20,680 in materials cost and $17,320 in conversion cost. The June 1 work in process inventory was 100% complete with respect to materials and 60% complete with respect to conversion. During June, 37,800 units were started into production. The June 30 work in process inventory consisted of 8,600 units that were 100% complete with respect to materials and 50% complete with respect to conversion.

1-1. How many units were completed and transferred to finished goods during the period?

1-2. Compute the equivalent units of production for materials.

1-3. Compute the equivalent units of production for conversion.

2-1. What is the cost of beginning work in process inventory plus the cost added during the period for materials?

2-2. What is the cost of beginning work in process inventory plus the cost added during the period for conversion?

3. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.)
Answer #1

1-1) Unit completed and transferred out = 5300+37800-8600 = 34500 Units

1-2) Equivalent unit of material = 34500+8600 = 43100

1-3) Equivalent unit of conversion = 34500+(8600*50%) = 38800

2-1) Total Cost of material = 20680+123705 = 144385

2-2) Total Cost of conversion = 199820

3) Cost per equivalent unit for material = 144385/43100 = 3.35